hartnn (hartnn):

Anyone getting bored? try this : If a³ + b³ = 0 , then the value of log (a + b) – 0.5(log a + log b + log 3) is equal to ??

hartnn (hartnn):

just for safety: a\(\ne\)b a\(\ne\)-b

Parth (parthkohli):

\[a^3 + b^3 = (a + b)(a^2 - ab + b^2)\]So,\[(a + b)(a^2 - ab + b^2) = 0\]Dividing both sides by a + b,\[a^2 - ab + b^ 2 = 0\]Since \(a^2 - 2ab + b^2 = (a - b)^2\), we have,\[a^2 - 2ab + b^2 +ab = 0\]\[(a - b)^2 + ab = 0\]\[(a - b)^2 = - ab\]\[a^2 + b^2 = ab\]One such case is a = 0 and b = 0, but you haven't allowed that. :(

hartnn (hartnn):

hmm....nice try no other formula for a^3+b^3 ?

OpenStudy (nipunmalhotra93):

Are you sure that your question is correct? If a^3 + b^3=0, then either a=-b or a=b=0.... :\ (because x^3 if a one one function)

hartnn (hartnn):

yes, question is correct.

OpenStudy (nipunmalhotra93):

0

hartnn (hartnn):

how ?

OpenStudy (nipunmalhotra93):

continuing from parthkohli,